# A Reinforcement Learning Approach to Synthetic Data Generation

### v1

@@ Abstract
Synthetic data generation (SDG) is a promising approach for enabling data sharing in biomedical studies while preserving patient privacy. Yet, state-of-the-art generative models often require large datasets and complex training procedures, limiting their applicability in small-sample settings. In this work, we reframe SDG as a reinforcement learning (RL) problem and introduce RLSyn, a novel framework that models the data generator as a stochastic policy over patient records and optimizes it using Proximal Policy Optimization with discriminator-derived rewards, yielding more stable and data-efficient training. We evaluate RLSyn on two biomedical datasets - AI-READI and MIMIC-IV - and benchmark it against state-of-the-art generative adversarial networks (GANs) and diffusion-based methods across extensive privacy, utility, and fidelity evaluations. RLSyn performs comparably to diffusion models and outperforms GANs on MIMIC-IV, while outperforming both diffusion models and GANs on the smaller AI-READI dataset. These results demonstrate that reinforcement learning provides a principled and effective alternative for synthetic biomedical data generation, particularly in data-scarce regimes.
@@ 1 Introduction
The amount of biomedical data that is collected through healthcare and research continues to expand at a rapid rate. Unfortunately, our ability to share such data remains limited due to numerous challenges, including concerns over the privacy of patients and research participants [kalkman_patients_2022, schreiber_what_2024, olsen_worldwide_2025], risk aversion [varhol_attitudes_2025, hermansen_heuristics_2023, stieglitz_when_2020], and general uncertainty over regulatory requirements at the local, national, and international [lencucha_trust_2021, gudi_regulatory_2022] level. One emerging approach to help mitigate a lack of data sharing is the notion of synthetic data [giuffre_harnessing_2023, extance_ai-generated_2025]. Modern synthetic data generation (SDG) methods are typically based on generative artificial intelligence (AI) that is trained to create new patient records that emulate the patterns of the original data, thereby preserving key biomedical relationships within the data while severing the link between real individuals [yan_generating_2024, noauthor_synthetic_2025].
Despite the promise of generative AI for SDG, state-of-the-art SDG methods have relied on the use of ever larger datasets, which, in turn, have facilitated the use of increasingly complex generative models [kaplan_scaling_2020]. As a result, current state-of-the-art approaches to SDG often leverage datasets with hundreds of thousands of data points [theodorou_synthesize_2023, yan_multifaceted_2022], limiting their applicability for the smaller datasets that are often used in biomedicine. This trend is driven, in part, by the growing dominance of diffusion-based generative models, which require a computationally demanding denoising process [ho_denoising_2020] that may be infeasible for small datasets. Additionally, prior approaches based on generative adversarial networks (GANs) rely on adversarial dynamics that are often unstable, data-inefficient, and prone to well-known failures, such as mode collapse.
@@
To test this hypothesis, we introduce RLSyn, an RL-based solution to SDG. We then compare RLSyn to existing GAN-based methods and state-of-the-art diffusion models on two publicly-available medical datasets. We conduct extensive experiments, assessing the utility, fidelity, and privacy of the synthetic data generated by these models. We find that RLSyn performs comparably to state-of-the-art SDG methods on larger datasets, while outperforming them on smaller datasets, highlighting the utility of RLSyn for SDG.
@@ 1.1 Statement of Significance
Who Would Benefit: Scientists seeking to share synthetic data for transparent and reproducible science whose work relies on small, curated datasets.
@@ 2 Related Work
We situate our work within prior research on synthetic data generation, focusing on generative adversarial networks and diffusion-based approaches. We then provide an overview of reinforcement learning and discuss its prior applications.
@@ 2.1 Generative Adversarial Networks
GANs are a class of deep learning models that learn to generate realistic data through an adversarial process between two neural networks- a generator that produces synthetic samples and a discriminator that attempts to distinguish between real and fake data [goodfellow_generative_2014]. Through this competition, the generator learns to produce data that closely approximates the true distribution of the input data. The use of GANs for SDG in healthcare was first identified by Choi et al. with the invention of medGAN [choi_generating_2017]. Since then, GANs have been studied for their ability to generate many different biomedical modalities, including imaging [jackson_enhancement_2024, frid-adar_gan-based_2018] and text in the form of clinical notes [baumel_controllable_2024, li_are_2021]. Additionally, much work has gone into improving GANs for tabular data generation [yan_multifaceted_2022, zhang_ensuring_2020, baowaly_synthesizing_2019]. The primary drawback of GANs is that it can be difficult to learn an effective discriminator. For example, if the discriminator is too strong and easily distinguishes real from synthetic samples, then the gradient updates become negligible and the generator does not improve. On the other hand, if the discriminator is too weak (i.e., it cannot effectively distinguish real from synthetic data), then it will fail to provide meaningful feedback to the generator [arjovsky_towards_2017]. These issues have been partially addressed by modifying the GAN loss function [arjovsky_wasserstein_2017]; however, instabilities in GAN training persist due to their sensitivity to hyperparameters and the non-stationarity of the loss function [lin_why_2021].
@@ 2.2 Denoising Diffusion Models
Denoising diffusion models [ho_denoising_2020] are a class of generative models that learn to synthesize data by reversing a gradual noising process - starting from random noise and iteratively denoising it to produce realistic samples. Diffusion models are typically implemented by adding noise to the original data and then training neural networks to remove that noise, reconstructing the original image. These models have exhibited state-of-the-art performance in generating high-quality synthetic images [ho_denoising_2020, rombach_high-resolution_2022]. For tabular healthcare data, several diffusion-based approaches have been developed, appearing to outperform their GAN-based counterparts [yuan_ehrdiff_2024, yoon_ehr-safe_2023]. While diffusion models are considered more principled approaches to SDG than GANs, they suffer from two non-trivial deficiencies in the context of SDG for healthcare. First, diffusion models have been shown to memorize training data, occasionally reproducing nearly identical copies, which may limit their applications to data sharing in healthcare [carlini_extracting_2023, rahman_investigating_2025]. Second, diffusion models require large amounts of training data, which may make them difficult to train from scratch for healthcare applications.
@@ 3.1 Datasets
The first dataset is based on AI-READI version 2.0, which contains data on 1067 patients spanning the full spectrum of type 2 diabetes severity, collected across multiple sites in the United States between 2023 and 2024. We extracted tabular data on participant diagnoses and laboratory measurements, as well as automatic measurements from electrocardiograms. Additionally, participants in AI-READI spent several days each with wearable activity monitors and continuous glucose monitors. We processed this time-series data by temporally aligning measurements from different sensors onto a common time grid and interpolating short gaps arising from irregular sampling or missing values. Days with insufficient sensor coverage or extended gaps that could not be reliably interpolated were excluded. We then split the remaining data by person-days and computed the number of hours in each of the four sleep stages and the four activity stages. We computed the mean heart rate and respiratory rate over the course of the day as well as the mean and standard deviation of their blood glucose. We computed participants’ resting heart rate during sedentary periods and their total calories burned and steps over each day. Lastly, we computed the mean of a stress score provided by the fitness wearable. We concatenated the 16 wearable-derived features with 56 continuous features consisting of clinical measurements (e.g., insulin levels) and measures derived from electrocardiograms (e.g., QT duration). We combined the resulting 72 continuous features with the most commonly-recorded 36 clinical conditions, yielding a dataset of 4,537 person-days (rows) of tabular data, each with 108 features.
@@ 3.2 RLSyn
In practice, the policy distribution is implemented as a combination of continuous and categorical components. For continuous features, the generator outputs the mean and variance that parameterize a Gaussian distribution, from which samples are drawn, passed through a $\tanh$ transformation and constrained to the range $(0,1)$. For categorical and binary features, the generator outputs logits that parameterize independent Bernoulli distributions. Together, these define a joint distribution $p_{\theta}(x)$ over all feature values, representing the generator’s probabilistic policy. The generator also maintains a value baseline that is relied upon during policy optimization. Figure 1 illustrates the generator architecture.
@@ 3.3 Baselines
We compared RLSyn against two baseline generative models developed specifically to promote privacy in the sharing of tabular health data: Electronic Medical Record Wasserstein GAN (EMR-WGAN) [zhang_ensuring_2020] and Electronic Health Record Diffusion (EHRDiff) [yuan_ehrdiff_2024]. EMR-WGAN was specifically designed for its use in structured EHR data and has shown strong performance across multiple datasets, outperforming other GAN architectures [yan_multifaceted_2022]. EHRDiff is a diffusion model that, like EMR-WGAN, was developed specifically for structured EHR data. It has been shown to outperform EMR-WGAN on EHR generation, while maintaining privacy. Collectively, these models represent existing state-of-the-art algorithms for tabular health record generation.
@@ 3.4.3 Fidelity
Similar to the utility evaluations, we trained models on the real training data and tested them on the synthetic data, a setting we refer to as real-to-synthetic or R2S. This analysis captures two main components of data fidelity. First, as with S2R, if classification performance matches that of the real data, this indicates that the same statistical relationships in the real data are present in the synthetic data. Second, the R2S analysis trains the classifier on the same real data that the synthetic data was generated from. Consequently, if the R2S results are higher than the R2R baseline, this indicates that the generative model may have simplified the synthetic data by removing outliers, over-fitting, or replicating close proxies of the real data.
@@ 3.5 Model Training
We partitioned AI-READI and MIMIC-IV into 90-10 and 70-30 training and testing datasets, respectively. To ensure the robustness of our analysis, we tuned the hyperparameters of each SDG model-dataset pair for 20 trials using Optuna [10.1145/3292500.3330701], selecting the hyperparameters that maximized classification performance for the S2R utility analysis. The final hyperparameters chosen are provided in Appendix Table A.3. After selecting the optimal hyperparameters, we trained each SDG model 10 times for each dataset across 10 different train-test splits and averaged all performance measures accordingly.
@@ 4.2 Utility
Figure 2, row 1 summarizes the the utility results. The S2R analysis on the MIMIC-IV dataset indicated that EHRDiff and RLSyn were the best-performing with respectable classification AUCs of 0.906 and 0.902, respectively. EMR-WGAN trailed with an AUC of 0.764. Similar results were observed in the AI-READI dataset. Specifically, RLSyn and EHRDiff had comparable S2R performance (AUC: 0.873 and 0.871, respectively), both outperforming EMR-WGAN (AUC: 0.747).
@@ 4.3 Fidelity
For the latent space comparison, clustering was performed using the elbow method, resulting in 3 clusters for AI-READI and 8 clusters for MIMIC-IV. The results for the other fidelity metrics mirror those of the R2S analysis in that RLSyn had the lowest NMI between clusters and CWC difference, outperforming both EMR-WGAN and EHRDiff for AI-READI (Table 2). Similarly, for MIMIC-IV RLSyn and EHRDiff were similar in terms of NMI (0.001 vs 0.003) and CWC (8.877 vs 7.848). Notably, both sets of values were much lower than their counterparts generated by EMR-WGAN (NMI: 0.320, CWC: 11.852). These differences are made visually apparent in Figure 3, which plots the principal components of the real and synthetic datasets. It can be seen that that RLSyn much more closely matches the real data distribution.
@@ 5 Discussion
Our findings indicate that EHRDiff and RLSyn outperformed EMR-WGAN in terms of utility and fidelity on both the MIMIC-IV and AI-READI datasets. Performance between RLSyn and EHRDiff was more competitive. On MIMIC-IV, RLSyn and EHRDiff achieved comparable utility and privacy, while RLSyn outperformed EHRDiff in terms of fidelity. However, on AI-READI, while RLSyn and EHRDiff still achieved comparable utility, RLSyn outperformed EHRDiff in terms of fidelity. Additionally, RLSyn achieved this performance with greater privacy than EHRDiff. These results suggest that RLSyn is more data-efficient than both GANs and diffusion approaches, producing higher-quality synthetic data with minimal privacy risks.
@@ Appendix A Appendix
Note: No tuning was performed for EMR-WGAN on MIMIC-IV as we used the exact hyperparameters and code as described in [yan_generating_2024].

### v2

@@ Abstract
Objective: Synthetic data generation (SDG) is a promising approach for enabling data sharing in biomedical studies while preserving patient privacy. Yet, state-of-the-art generative models often require large datasets and complex training procedures, limiting their applicability in small-sample settings common in biomedical research. This study aims to develop a more principled and efficient approach to SDG and evaluate its efficacy for biomedical applications.
Methods: In this work, we reframe SDG as a reinforcement learning (RL) problem and introduce RLSyn, a novel framework that models the data generator as a stochastic policy over patient records and optimizes it using Proximal Policy Optimization with discriminator-derived rewards. We evaluate RLSyn on two biomedical datasets–AI-READI and MIMIC-IV–and benchmark it against state-of-the-art generative adversarial networks (GANs) and diffusion-based methods across extensive privacy, utility, and fidelity evaluations.
Results: On MIMIC-IV, RLSyn achieves predictive utility comparable to diffusion models (S2R AUC 0.902 vs 0.906 respectively) while slightly outperforming them in fidelity (NMI 0.001 vs. 0.003; DWD 2.073 vs. 2.797) and achieving comparable, low privacy risk ( 0.50 membership inference risk AUC). On the smaller AI-READI dataset, RLSyn again matches diffusion-based utility (S2R AUC 0.873 vs. 0.871), while achieving higher fidelity (NMI 0.001 vs. 0.002; DWD 13.352 vs. 16.441) and significantly lower vulnerability to membership inference attacks (AUC 0.544 vs. 0.601). Both RLSyn and diffusion-based models substantially outperform GANs across utility and fidelity on both datasets.
Conclusion: Reinforcement learning provides a principled and effective approach for synthetic biomedical data generation, particularly in data-scarce regimes.
@@ 1 Introduction
The amount of biomedical data that is collected through healthcare and research continues to expand at a rapid rate. Unfortunately, our ability to share such data remains limited due to numerous challenges, including concerns over the privacy of patients and research participants [kalkman_patients_2022, schreiber_what_2024, olsen_worldwide_2025], risk aversion [varhol_attitudes_2025, hermansen_heuristics_2023, stieglitz_when_2020], and general uncertainty regarding regulatory requirements at the local, national, and international level [lencucha_trust_2021, gudi_regulatory_2022] . One emerging approach for mitigating these barriers to data sharing is the use of synthetic data. [giuffre_harnessing_2023, extance_ai-generated_2025]. Modern synthetic data generation (SDG) methods are typically based on generative artificial intelligence (AI) that is trained to create new patient records that emulate the patterns of the original data, thereby preserving key biomedical relationships within the data while severing the link to real individuals [yan_generating_2024, noauthor_synthetic_2025].
Despite the promise of generative AI for SDG, state-of-the-art SDG methods have relied on the use of ever larger datasets, which, in turn, have facilitated the use of increasingly complex generative models [kaplan_scaling_2020]. As a result, current state-of-the-art approaches to SDG often leverage datasets with hundreds of thousands of data points [theodorou_synthesize_2023, yan_multifaceted_2022], limiting their applicability for the smaller datasets that are often used in biomedical research. This trend is driven, in part, by the growing dominance of diffusion-based generative models, which require a computationally demanding denoising process [ho_denoising_2020] that may be infeasible for small datasets. Additionally, prior approaches based on generative adversarial networks (GANs) rely on adversarial dynamics that are often unstable, data-inefficient, and prone to well-known failures, such as mode collapse.
@@
To test this hypothesis, we introduce RLSyn, an RL-based framework for SDG. We then compare RLSyn to existing GAN-based methods and state-of-the-art diffusion models on two publicly-available medical datasets. We conduct extensive experiments, assessing the utility, fidelity, and privacy of the synthetic data generated by these models. We find that RLSyn performs comparably to state-of-the-art SDG methods on larger datasets, while outperforming them on smaller datasets, highlighting the utility of RLSyn for SDG.
@@ 1.1 Statement of Significance
Who Would Benefit: Scientists seeking to share synthetic data for transparent and reproducible science whose work is based on small, curated datasets.
@@ 2 Related Work
We situate our work within prior research on SDG, focusing on generative adversarial networks and diffusion-based approaches. We then provide an overview of reinforcement learning and discuss its prior applications.
@@ 2.1 Generative Adversarial Networks
GANs are a class of deep learning models that learn to generate realistic data through an adversarial process between two neural networks: a generator that produces synthetic samples and a discriminator that attempts to distinguish between real and fake data [goodfellow_generative_2014]. Through this competition, the generator learns to produce data that closely approximates the true distribution of the input data. The use of GANs for SDG in healthcare was first identified by Choi et al. with the invention of medGAN [choi_generating_2017]. Since then, GANs have been studied for their ability to generate many different biomedical modalities, including imaging [jackson_enhancement_2024, frid-adar_gan-based_2018] and text in the form of clinical notes [baumel_controllable_2024, li_are_2021]. Additionally, much work has gone into improving GANs for tabular data generation [yan_multifaceted_2022, zhang_ensuring_2020, baowaly_synthesizing_2019]. The primary drawback of GANs is that it can be difficult to learn an effective discriminator. For example, if the discriminator is too strong and easily distinguishes real from synthetic samples, then the gradient updates become negligible and the generator does not improve. On the other hand, if the discriminator is too weak (i.e., it cannot effectively distinguish real from synthetic data), then it will fail to provide meaningful feedback to the generator [arjovsky_towards_2017]. These issues have been partially addressed by modifying the GAN loss function [arjovsky_wasserstein_2017]; however, instabilities in GAN training persist due to their sensitivity to hyperparameters and the non-stationarity of the loss function [lin_why_2021].
@@ 2.2 Denoising Diffusion Models
Denoising diffusion models [ho_denoising_2020] are a class of generative models that synthesize data by learning to reverse a gradual noising process, in which noise is incrementally added to real data and a neural network is trained to predict and remove this noise at each step. At generation time, the model starts from random noise and iteratively denoises it to produce realistic samples. These models have exhibited state-of-the-art performance in generating high-quality synthetic images [ho_denoising_2020, rombach_high-resolution_2022]. For tabular healthcare data, several diffusion-based approaches have been developed, appearing to outperform their GAN-based counterparts [yuan_ehrdiff_2024, yoon_ehr-safe_2023]. While diffusion models are considered more principled approaches to SDG than GANs, they suffer from two non-trivial deficiencies in the context of SDG for healthcare. First, diffusion models have been shown to memorize training data, occasionally reproducing nearly identical copies, which may limit their applications to data sharing in healthcare as this may pose a risk to data privacy [carlini_extracting_2023, rahman_investigating_2025]. Second, diffusion models require large amounts of training data, which may make them difficult to train from scratch for healthcare applications.
@@ 3.1 Datasets
The first dataset is based on AI-READI version 2.0, which contains data on 1067 patients spanning the full spectrum of type 2 diabetes severity, collected across multiple sites in the United States between 2023 and 2024. We extracted tabular data on participant diagnoses and laboratory measurements, as well as automatic measurements from electrocardiograms. Additionally, participants in AI-READI spent several days wearing wearable activity monitors and continuous glucose monitors. We processed the wearable time-series data by aligning measurements from multiple sensors onto a common daily time grid with 5-minute resolution, anchored at 08:00 AM. Short gaps due to irregular sampling were interpolated, with interpolation restricted to gaps of at most three hours. Days with insufficient sensor coverage, extended gaps beyond this threshold, or physiologically implausible values (e.g., calorie expenditure exceeding 20 kcal/min) were excluded. We then split the remaining data by person-days and computed the number of hours in each of the four sleep stages and the four activity stages. We computed the mean heart rate and respiratory rate over the course of the day as well as the mean and standard deviation of their blood glucose. We computed participants’ resting heart rate during sedentary periods and their total calories burned and steps over each day. Lastly, we computed the mean of a stress score provided by the fitness wearable. We concatenated the 16 wearable-derived features with 56 continuous features consisting of clinical measurements (e.g., insulin levels) and measures derived from electrocardiograms (e.g., QT duration). We combined the resulting 72 continuous features with the most commonly-recorded 36 clinical conditions, yielding a dataset of 4,537 person-days (rows) of tabular data, each with 108 features. A complete list of features can be found in Appendix Table 4.
@@ 3.2 RLSyn
In practice, the policy distribution is implemented as a combination of continuous and categorical components. For continuous features, the generator outputs the mean and variance that parameterize a Gaussian distribution, from which samples are drawn, passed through a $\tanh$ transformation and constrained to the range $(0,1)$. For categorical and binary features, the generator outputs logits that parameterize independent Bernoulli distributions. Together, these define a joint distribution $p_{\theta}(x)$ over all feature values, representing the generator’s probabilistic policy. The generator also maintains a value baseline that is utilized during policy optimization. Figure 1 illustrates the generator architecture.
@@ 3.3 Baselines
We compared RLSyn against two baseline generative models developed specifically to promote privacy in the sharing of tabular health data: Electronic Medical Record Wasserstein GAN (EMR-WGAN) [zhang_ensuring_2020] and Electronic Health Record Diffusion (EHRDiff) [yuan_ehrdiff_2024]. EMR-WGAN was specifically designed for its use in structured EHR data and has shown strong performance across multiple datasets, outperforming other GAN architectures [yan_multifaceted_2022]. EHRDiff is a diffusion model that, like EMR-WGAN, was developed specifically for structured EHR data. It has been shown to outperform EMR-WGAN on specific EHR data generation tasks, while maintaining privacy. Collectively, these models represent existing state-of-the-art algorithms for tabular health record generation.
@@ 3.4.3 Fidelity
Similar to the utility evaluations, we trained models on the real training data and tested them on the synthetic data, a setting we refer to as real-to-synthetic or R2S. As with S2R, if R2S classification performance matches that of models trained and tested on the real data, this indicates that the same statistical relationships in the real data are present in the synthetic data. While some prior work has considered R2S to be a utility metric, we classify it as a fidelity measure. This is because R2S values that are higher than performance on the real data do not indicate higher quality synthetic data. A key detail here is that the R2S analysis trains the classifier on the same real data that the synthetic data was generated from. Consequently, if the R2S results are higher than the R2R baseline, this indicates that the generative model may have simplified the synthetic data by removing outliers, over-fitting, or replicating close proxies of the real data. As a result we aim for S2R classification performance that matches, but does not exceed the real data counterpart.
@@ 3.5 Model Training
We partitioned AI-READI and MIMIC-IV into 90-10 and 70-30 training and testing datasets, respectively. To ensure the robustness of our analysis, we tuned the hyperparameters of each SDG model-dataset pair for 20 trials using Optuna [10.1145/3292500.3330701], selecting the hyperparameters that maximized classification performance for the S2R utility analysis. However, for EMR-WGAN on MIMIC-IV, we did not perform additional tuning, as we used the exact code and hyperparameters as described in [yan_generating_2024]. The final hyperparameters chosen are provided in Appendix Table A.3. After selecting the optimal hyperparameters, we trained each SDG model 10 times for each dataset across 10 different train-test splits and averaged all performance measures accordingly.
@@ 4.2 Utility
Figure 2, row 1 summarizes the utility results. The S2R analysis on the MIMIC-IV dataset indicated that EHRDiff and RLSyn were the best-performing with respectable classification AUCs of 0.906 and 0.902, respectively. EMR-WGAN trailed with an AUC of 0.764. Similar results were observed in the AI-READI dataset. Specifically, RLSyn and EHRDiff had comparable S2R performance (AUC: 0.873 and 0.871, respectively), both outperforming EMR-WGAN (AUC: 0.747).
@@ 4.3 Fidelity
For the latent space comparison, clustering was performed using the elbow method, resulting in 3 clusters for AI-READI and 8 clusters for MIMIC-IV. The results for the other fidelity metrics mirror those of the R2S analysis in that RLSyn had the lowest NMI between clusters and CWC difference, outperforming both EMR-WGAN and EHRDiff for AI-READI (Table 2). Similarly, for MIMIC-IV RLSyn and EHRDiff were similar in terms of NMI (0.001 vs 0.003) and CWC (8.877 vs 7.848). Notably, both sets of values were much lower than their counterparts generated by EMR-WGAN (NMI: 0.320, CWC: 11.852). These differences are made visually apparent in Figure 3, which plots the principal components of the real and synthetic datasets. It can be seen that that RLSyn more closely matches the real data distribution for both datasets.
@@ 5 Discussion
Our findings indicate that EHRDiff and RLSyn outperformed EMR-WGAN in terms of utility and fidelity on both the MIMIC-IV and AI-READI datasets. Performance between RLSyn and EHRDiff was more competitive. On MIMIC-IV, RLSyn and EHRDiff achieved comparable utility and privacy, while RLSyn outperformed EHRDiff in terms of fidelity. However, on AI-READI, while RLSyn and EHRDiff still achieved comparable utility, RLSyn outperformed EHRDiff in terms of fidelity and achieved this performance with greater privacy than EHRDiff. These results suggest that RLSyn is more data-efficient than both GANs and diffusion approaches, producing higher-quality synthetic data with minimal privacy risks.
@@ Appendix A Selected Hyperparameters
PPO Mean Penalty 0.0 N/A N/A 0.2 N/A N/A
PPO Clip 0.1 N/A N/A 0.1 N/A N/A
PPO Entropy Coef. 0.001 N/A N/A 0.001 N/A N/A
PPO Value Coef. 0.5 N/A N/A 0.5 N/A N/A
  • 1.
    Note. No tuning was performed for EMR-WGAN on MIMIC-IV; we used the hyperparameters and code from [yan_generating_2024]. For RLSyn, PPO clip, PPO entropy weight, and PPO value weight were not tuned, as these follow standard PPO defaults.
Appendix B AI-READI Features
Table 4: Complete list of AI-READI features used (108 total).
Wearable and glucose monitor-derived continuous features
Mean Heart Rate Mean Blood Glucose Mean Respiration Rate
Mean Stress Std Blood Glucose Resting Heart Rate
Total Steps Total Calories Burned Light Sleep Hours
Deep Sleep Hours REM Sleep Hours Awake Hours
Generic Activity Hours Walking Activity Hours Running Activity Hours
Sedentary Activity Hours
Clinical, cognitive, and ECG continuous features
A/G Ratio vaflue ALT (IU/L) AST (IU/L)
Albumin (g/dL) Alkaline Phosphatase (IU/L) BUN (mg/dL)
BUN/Creatinine ratio Bilirubin Total (mg/dL) C-Peptide (ng/mL)
CRP–HS (mg/L) Calcium (mg/dL) Carbon Dioxide Total (mEq/L)
Chloride (mEq/L) Creatinine (mg/dL) Globulin Total (g/dL)
Glucose (mg/dL) HDL Cholesterol (mg/dL) Insulin (ng/mL)
LDL Cholesterol (mg/dL) Potassium (mEq/L) Protein Total (g/dL)
Sodium (mEq/L) Total Cholesterol (mg/dL) Triglycerides (mg/dL)
Urine Albumin (mg/DL) Urine Creatinine (mg/DL) Digit Span Time value
Delayed Recall With No Clue Time value Clock Visuospatial Executive Time value Cube Visuospatial Executive Time value
Memory Trial 1 Time value Memory Trial 2 Time value Moca Abstraction Time value
Moca Orientation Time value Moca Total Score Time value Naming Time value
Subtraction Time value Repetition Time value Lettera Time value
Trails Visuospatial Executive Time value Moca Combined Mis Score value Memory Trial 1 value
Memory Trial 2 value Moca Abstraction value Moca Orientation value
Naming value Repetition value Subtraction value
PR (ECG) QRSD (ECG) QT (ECG)
QTc (ECG) P (ECG) QRS (ECG)
T (ECG)
Categorical condition and diagnostic features
Age-related macular degeneration Arthritis Cancer
Cataracts (1+ eyes) Chronic pulmonary problems Circulation problems
Diabetic retinopathy (1+) Digestive problems Marijuana user
Dry eye (1+) Glaucoma (1+) Hearing impairment
Heart attack High blood cholesterol High blood pressure
Kidney problems Low blood pressure Mild cognitive impairment
Multiple sclerosis Obesity Osteoporosis
Other heart issues (pacemaker) Other neurological conditions Parkinson’s disease
Retinal vascular occlusion Stroke Type 2 Diabetes
Urinary problems Abnormal (ECG) Borderline (ECG)
Normal (ECG) Otherwise Normal (ECG) Fluency Language value
Cube Visuospatial Executive value Lettera value Trails Visuospatial Executive value
A complete list of all AI-READI features used in our experiments is provided in Table 4.